OpenStudy (anonymous):

Ann and Jane bowl regularly at Strike 'b Spares Lanes. Last Tuesday Ann and Jane had a combined score of 425. Ann's score was (equaled) 25 more than Jan's score. Find their scores.

OpenStudy (nathanjhw):

200 and 225.

OpenStudy (texaschic101):

A + J = 425 A = J + 25 now we sub (J + 25) + J = 425 2J + 25 = 425 2J = 425 - 25 2J = 400 J = 400/2 J = 200 A + J = 425 A + 200 = 425 A = 425 - 200 A = 225 check.. A + J = 425 A = J + 25 225 + 200 = 425 225 = 200 + 25 425 = 425 (correct) 225 = 225 (correct) therefore, Ann (A) made a score of 225 and Jane (J) made a score of 200.... so that is how it is done...I wanted to show you so you would understand. If you have any questions, please ask :)